# Protocol Futuring: Speculating Second-Order Dynamics of Protocols in Sociotechnical Infrastructural Futures

### v1

@@ 3.1. Process Overview
In practical terms, a typical Protocol Futuring framework might follow these steps:
@@
    Choose a Focal Tensions: This could be a real emerging protocol (e.g., a next-generation social media standard, a climate treaty framework in negotiation) or a hypothetical one addressing a societal challenge (e.g., a protocol for managing AI “rights” or a protocol for inter-city resource sharing). Sometimes we frame it broadly (“imagine the future of education protocols”) to allow diverse sub-topics. The tension should be both technically plausible and socially consequential—e.g., authenticity vs interpretability, liability vs autonomy, privacy vs auditability. Effective tensions usually involve trade-offs that cannot be resolved solely through technical design.
@@
    Gather Research and Inspiration: Before jumping into fiction, we compile insights from current technology trends, history, and social context. For instance, if speculating about knowledge archive protocols, we look at current digital preservation efforts, historical library systems, and even myths about archives (to inspire creative angles). This grounding ensures the fiction has anchors in reality, making it more relatable and plausibly extrapolated.
@@
    Workshop or Creative Session: If participatory, convene a session (could be one day or multi-day) with stakeholders and creatives. Use exercises to explore the space—mapping stakeholders and their goals, identifying key tensions/trade-offs (privacy vs transparency, efficiency vs resilience, central control vs decentralization, etc.), and brainstorming potential futures. We often encourage teams to formulate a future scenario matrix: e.g., vary two factors to create four quadrants of futures, and then flesh out stories in each. In a contest format (like Ghosts in Machines), this workshop phase is decentralized—individual writers brainstorm on their own but are given prompts and guidelines to stimulate thinking around the chosen theme.
@@
    Protocol Creation: Participants produce the fictional protocols and narratives. We stress depicting concrete situations: e.g., a day in the life of someone governed by the protocol, a breaking news story about a protocol failure, a personal diary from a maintainer of the system, etc. These stories are where imaginative leaps happen, but within the constraints identified. The artifacts might be accompanied by short explanatory text if needed (as an academic project, we sometimes include a rationale section for each fiction that connects it back to research questions).
@@
    Speculation and Analysis: After creation, the research team analyzes the fictions for second-order dynamics insights. For example, do multiple independent stories all foresee a particular risk? Or, do they propose interesting design features for protocols (like kill-switches, consent mechanisms, fallback modes) that could be worth exploring?
@@
    Iterate or Refine: Early rounds often reveal blind spots—for example, scenarios that skew overly optimistic, overly technical, or insufficiently attentive to social conflict. In such cases, we may deliberately introduce a ”Blue Team” / ”Red Team” iteration: a Blue Team extends or stabilizes the existing protocol world, while a Red Team systematically challenges its assumptions, surfaces vulnerabilities, or explores adversarial interpretations. Because fiction accommodates multiple coexisting truths, iteration does not require selecting a single “correct” future. Instead, we refine the design space by adding counter-fictions, layering competing perspectives, or tweaking earlier contributions to expose alternative trajectories and tensions.
@@ 4. Case Study: Knowledge Futurama
To demonstrate the utility and versatility of Protocol Futuring, we focus in depth on a single flagship case: Knowledge Futurama, a multi-team participatory fiction exercise that most clearly embodies the framework introduced in the previous section. Knowledge Futurama allows us to trace, with ethnographic granularity, how protocol rules, mediating practices, and second-order effects emerge through iterative inheritance and reinterpretation—making it an ideal primary case for illustrating the method’s analytical contribution. A second case, South Beast Asia, is included in the appendix as a supplementary example. Whereas Knowledge Futurama centers on long-term knowledge preservation and highlights the multi-century drift of infrastructural protocols, South Beast Asia explores regional AI governance futures and demonstrates the method’s applicability to geopolitical and socio-cultural contexts. We include it to show range, but concentrate our analysis on Knowledge Futurama because it provides the clearest instantiation of Protocol Futuring’s epistemic logic.
4.1. Context and Goal of Knowledvge Futurama
Knowledge Futurama was a participatory futures workshop conducted as part of the Summer of Protocols 2025 program at the Edge Esmeralda gathering in California. Around thirty participants—including technologists, designers, librarians, artists, and futurists—explored how knowledge might be preserved and made meaningful across a millennium. The aim was to prototype “protocols for 1,000-year libraries”: sociotechnical arrangements that could keep knowledge alive, intelligible, and legitimate through centuries of political, ecological, and technological change. The name echoed the 1939 World’s Fair “Futurama” exhibit, but turned its speculative lens toward long-term memory infrastructures rather than consumer futures.
Two of the authors participated directly in the workshop as embedded researchers. A participant-observer stance (Dourish 2006) has been adopted, where we documented the evolving scenarios, mediating practices, and interpretive ambiguities as they unfolded. Following the event, we conducted elicitation interviews with two workshop participants (one of whom was an author of this article, indexed as “Interviewee 2”) to reconstruct design rationales, clarify interpretive moves, and surface participants’ own theories of preservation, drift, and relevance. These interviews were not intended to provide a representative sampling but to generate situated, experiential insight into how actors reasoned through long-term protocol design and how the relay structure shaped their thinking. We framed the core design challenge as building a preservation system that is not simply about storing data, but about maintaining usable protocols that can evolve. Participants emphasized a nearer design horizon of 100 years as a meaningful unit for institutional and infrastructural commitments, while also using speculative devices to explore what happens as protocols outlive their creators.
@@ 4.2. Process and Setting
The workshop was structured as a relay across six parallel story worlds. Participants were divided into tables of five and worked through six sequential rounds, with each table contributing to every scenario in turn. After the sixth round, each story is returned to its original authors. This allowed them to see how their initial protocol had drifted, mutated, or been reinterpreted by others, rendering interpretive drift visible and allowing them to experience how protocols evolve when traversed across many hands with only partial context.
@@
Each story world began with a “Blue Build 1” phase, where a team proposed a knowledge-preservation protocol for the next 100 years. This required teams to articulate both a technical stack (e.g., encoding methods, storage substrates, verification technologies) and a social milieu (e.g., maintenance institutions, stewarding communities, norms, governance practices). In the subsequent round, a new team inherited this design and introduced a “Red Crisis 1” intervention that represented political, ecological, social, or technological shocks that destabilized or reinterpreted the original protocol. The third round, Blue Build 2, tasked yet another team with continuing the story under these altered conditions for the next 300 years. They had to adapt, improvise, or recompose the preservation protocol in light of the disruptions and ambiguities left by their predecessors. This was followed by Red Crisis 2, where yet another group introduced further disruptions, forcing reinterpretations, repair, or workaround practices. This alternating pattern of “build” and crisis” repeats one last time with Blue Build 3 extending the protocol into the far future (the next 600 years), after which Red Crisis 3 introduced a last set of shocks, before the story was returned to the original authors. Over the course of this process, each knowledge artifact developed a layered, multi-century “protocol biography”: a record of how initial rules, assumptions, and incentives accumulated unintended consequences over centuries.
@@
The workshop was conducted entirely through six shared Google Doc files, one per story world, edited sequentially by all thirty participants. Each table deliberated in person while one designated note-taker captured their decisions around technology stacks, social milieus, crises, and state conditions directly into the evolving document. Because teams worked from what previous groups had written – brief bullet points, shorthand, half-formed sentences, and ambiguous phrasing – later teams frequently had to infer missing rationale, decide whether a line was meant as a firm rule or a tentative thought, and interpret text that reflected the stylistic quirks of an earlier note-taker rather than any stable consensus. No communication was allowed between tables until the final debrief, amplifying the interpretive gaps and making drift a constitutive part of the design process. This structure deliberately reproduced the dynamics of protocol evolution in the wild, where specifications, proposals, patches, and lived practices often sediment into infrastructures without any single authorial center or developmental trajectory. As in actual decentralized systems, protocols themselves were the design artifacts, rather than the fictional worlds, but the rules, mediations, and assumptions that structured their long-term behavior. The shared documents functioned as proto-infrastructures that were mutable, cumulative, and open for reinterpretation, reflecting how technical rules and social practices co-produce one another over time.
@@ 4.3.2. Red Crisis One (2120–2125): Authenticity Under Siege
4.3.3. Blue Build Two and Red Crisis Two (2125–2625): Hash Wars and Environmental Entanglement
Subsequent teams attempted to repair the jam by introducing new layers of protocol: comet-based “carbon dating” of engraved replicas, sandboxed LLMs to reconstruct pre-LLM word distributions, and a blockchain-verified hash of the “oldest known Wikipedia.” Interviewed participants described this as an attempt to build a “protocol hardness”– a durability that seeks to anchor truth in mechanisms that could not easily be spoofed or reinterpreted. At the same time, the social milieu shifted. AI researchers, amateur astronomy clubs, and archivists formed a coalition around maintaining a “clean” pre-LLM dataset. This coalition reinforced the idea that some slices of knowledge must remain stable referents, even as everything else is up for reinterpretation.
@@ 4.3.3. Blue Build Two and Red Crisis Two (2125–2625): Hash Wars and Environmental Entanglement
By 2625, distorted information had become culturally normalized, while undistorted data was derided as naïve or heretical. In this stage, a protocol designed to verify authenticity through intensive hash comparison and space-based signaling became entangled with heavy compute demands, regulatory oversight, and adversarial exploitation, ultimately producing second-order effects such as environmental externalities, political clampdowns, and a cultural turn that celebrated distortion over fidelity. The interviews helped clarify this trajectory as a risk-aware framing of long-term preservation: participants explicitly discussed political, economic, social, legal, and environmental constraints as integral to protocol design. Crisis recovery was not just about restoring systems, but about maintaining legitimacy under shifting constraints.
@@ 4.4. Design Principles and Mediating Practices
Third, participants treated crisis and recovery as intrinsic to protocol design. Several emphasized the need for “recovery protocols” capable of restoring function after disruption without sacrificing interpretability. This anticipatory stance toward failure differentiates Protocol Futuring from more static forms of design fiction, which often portray artifacts as working as intended within a closed narrative world. Here, resilience, repair, and flexible adaptation were core design considerations. Fourth, participants explicitly noted how philosophical and cultural attitudes—such as skepticism toward historical truth or tendencies toward artistic reinterpretation—might shape the long-term fate of preservation systems. Rather than assuming fidelity as a universal good, they saw value in designing infrastructures that honor historical authenticity while permitting creative reimagination. This orientation positions Protocol Futuring as a way to explore how future publics might negotiate the tension between accuracy and meaning. Finally, challenges around collaboration and communication surfaced repeatedly through metaphors of “different islands.” Because teams worked asynchronously through textual handovers, misinterpretation and semantic drift were routine features of the exercise. Participants proposed concepts such as “lingua defense”—shared languages, standards, and interpretive conventions—as cultural protocols to mitigate these breakdowns. This insight reframes communication and consensus mechanisms as preservation infrastructures in their own right, revealing how much interpretive labor underpins even seemingly technical systems.
@@ 5.1. Temporality, Second-Order Dynamics, and Protocol “Memory”
Second-order effects are inherently temporal: they emerge from how protocols are adopted, interpreted, and modified over time, often across generations. Protocol Futuring treats temporality not as a narrative backdrop but as a mechanism for making protocol behavior visible. Second-order effects are inherently temporal phenomena that unfold through repeated reinterpretation and sedimentation of rules. Traditional design futures often compress time into a single speculative moment, while infrastructural analyses typically examine systems synchronically. In contrast, Protocol Futuring operationalizes time as a structured sequence of inheritances, ruptures, and recoveries. By distributing protocol interpretation across multiple “future centuries” and multiple actors, PF makes temporal dynamics—such as eventual divergence, cyclical reinterpretation, or value decay—empirically accessible within a single session. This approach aligns with STS work on socio-technical imaginaries but renders long temporal arcs enacted rather than only conceptual. Our case study on Knowledge Futurama particularly highlights how Protocol Futuring can surface these temporal dynamics in ways that resonate with STS work on infrastructure time and sociotechnical imaginaries.
@@
In Knowledge Futurama, participants collectively designed archival protocols intended to persist for centuries. The resulting fictions centered not only on immediate trade-offs (e.g., privacy vs. access, centralization vs. federation) but also on how those choices sediment into long-term trajectories: protocols accumulating layers of patches, losing legibility, or being repurposed for unexpected ends. This echoes Ribes2007Tensions’s “long now” of infrastructure, where tensions between flexibility and stability, project timescales and infrastructural timescales, become central. Our speculative archive worlds dramatized, for instance, how a global archive protocol might gradually drift away from its founding rationale as institutional memory erodes and new stakeholders reinterpret its rules.
@@
Seen through STS work on temporality, these stories can be understood as miniature sociotechnical imaginaries attached to particular protocol regimes. jasanoff2015dreamscapes argue that large-scale projects—from nuclear power to smart grids—are guided by collectively held imaginaries of national futures. In our method, participants effectively articulate competing imaginaries of protocolized futures: worlds where archival AIs act as guardians of planetary memory versus worlds where protocol-enforced scarcity creates exclusion and information black markets. Protocol Futuring thus surfaces how different imaginaries are encoded into protocol choices, and how those imaginaries, in turn, shape what second-order phenomena are even thinkable.
@@
Our scenarios also foreground protocol “memory”: the ways in which protocols remember (or fail to remember) the world that produced them. Bowker’s work on biodiversity databases shows how infrastructures embody particular classificatory decisions that persist long after their original context has faded (bowker2000biodiversity). Similarly, the archive stories exhibit situations where only a small priesthood remembers why certain safeguards exist by the year 2500. Participants proposed design responses—such as embedding narrative origin stories or mythologies into code comments, protocol documents, or ritual practices—that explicitly tie technical rules to their founding controversies. This points toward protocol design strategies that treat documentation, storytelling, and education as first-class components of long-term governance, not mere afterthoughts.
@@
Finally, Protocol Futuring makes visible different temporal failure modes of protocols. Through our fictions, we can distinguish: (a) drift, where practice diverges from formal rules; (b) jam, where incompatible incentives or overloaded dependencies lead to contention and deadlock; and (c) ossification, where the cost or risk of change becomes so high that even beneficial modifications are repeatedly deferred. Each of these second-order dynamics maps onto observable phenomena in real infrastructures (e.g., workarounds and shadow systems, prolonged standardization disputes, or entrenched legacy protocols). Future empirical work can focus on observables such as divergence between protocol text and workaround artifacts (drift), frequency and distribution of change proposals (ossification), or patterns of stakeholder contestation (jam).

### v2

@@ 3.1. Process Overview
In practical terms, a typical Protocol Futuring engagement follows six steps:
@@
    Choose a Focal Tension: The process begins by identifying a sociotechnical tension that is both technically plausible and socially consequential. This may be an emerging real-world challenge (e.g., a next-generation social media standard or climate treaty framework) or a hypothetical scenario addressing competing stakeholder interests. Effective focal tensions involve trade-offs that cannot be resolved through technical design alone—such as authenticity versus interpretability, liability versus autonomy, or privacy versus auditability. The framing can be specific or broad (e.g., “imagine the future of education protocols”) depending on the desired scope. The protocol, conceived as an engineered argument, then becomes a vehicle for navigating these tensions.
@@
    Gather Research and Inspiration: Before engaging in fiction, participants compile background research on relevant technology trends, historical precedents, and social contexts. For example, speculation about knowledge archive protocols might draw on current digital preservation efforts, historical library systems, and cultural narratives about archives. This grounding anchors the fiction in reality, ensuring that speculative extrapolations remain relatable and plausible.
@@
    Workshop or Creative Session: Participatory engagements typically involve a structured session—often half a day—that brings together stakeholders and creative practitioners. Facilitation exercises guide participants through stakeholder mapping, tension identification (e.g., privacy versus transparency, centralization versus decentralization), and future brainstorming. A common technique is the scenario matrix: participants vary two key factors to generate four quadrant futures, then develop narratives within each quadrant.
@@
    Protocol Creation: Participants develop fictional protocols by articulating their rules and embedding them within narrative contexts. We encourage depicting concrete lifeworld situations—a day in the life of someone governed by the protocol, a news report about a protocol failure, or a diary entry from a system maintainer. These situated narratives enable imaginative exploration within the constraints established by the protocol rules. Artifacts may include explanatory annotations as needed.
@@
    Speculation and Analysis: Following creation, participants analyze the fictions to surface second-order dynamics. This analysis examines whether independent narratives converge on particular risks, or whether they propose design features worth further exploration (e.g., kill-switches, consent mechanisms, fallback modes). In participatory settings, this analytical work occurs in real time as participants inherit and respond to prior teams’ designs—interpreting ambiguous handovers, inferring unstated rationales, and deciding whether to extend or contest emerging protocol worlds.
@@
    Iterate and Refine: Initial rounds often reveal blind spots, such as scenarios that are overly optimistic, excessively technical, or insufficiently attentive to social conflict. To address this, we introduce adversarial iteration: a “Blue Team” extends or stabilizes the existing protocol world, while a “Red Team” challenges assumptions, surfaces vulnerabilities, and explores adversarial interpretations. Because fiction accommodates multiple coexisting truths, iteration does not require converging on a single “correct” future. Instead, the design space is refined through counter-fictions, competing perspectives, and revisions that expose alternative trajectories.
@@ 4. Case Study: Knowledge Futurama
To demonstrate the utility and versatility of Protocol Futuring, we focus in depth on a flagship case: Knowledge Futurama, a multi-team participatory fiction exercise that most clearly embodies the framework introduced in the previous section. Knowledge Futurama allows us to trace, with analytic detail, how protocol rules, mediating practices, and second-order effects emerge through iterative inheritance and reinterpretation—making it an ideal primary case for illustrating the method’s analytical contribution. Knowledge Futurama centers on long-term knowledge preservation and highlights the multi-century drift of infrastructural protocols.
4.1. Context and Goal of Knowledge Futurama
Knowledge Futurama was a participatory futures workshop conducted as part of the Summer of Protocols 2025  [^2] https://summerofprotocols.com/ program at the Edge Esmeralda  [^3] https://www.edgeesmeralda.com/ gathering in California. Around thirty participants—including technologists, designers, librarians, artists, and futurists—explored how knowledge might be preserved and made meaningful across a millennium. The aim was to prototype “protocols for 1,000-year libraries”: sociotechnical arrangements that could keep knowledge alive, intelligible, and legitimate through centuries of political, ecological, and technological change. The name echoed the 1939 World’s Fair “Futurama” exhibit, but turned its speculative lens toward long-term memory infrastructures rather than consumer futures.
Two of the authors participated directly in the workshop as embedded researchers. We adopted a participant-observer stance (Dourish, 2006), documenting the evolving scenarios, mediating practices, and interpretive ambiguities as they unfolded. Following the event, we conducted elicitation interviews with two workshop participants (one of whom was an author of this article, indexed as “Interviewee 2”) to reconstruct design rationales, clarify interpretive moves, and surface participants’ own theories of preservation, drift, and relevance. These interviews were not intended to provide a representative sampling but to generate situated, experiential insight into how actors reasoned through long-term protocol design and how the relay structure shaped their thinking. We framed the core design challenge as building a preservation system that is not simply about storing data, but about maintaining usable protocols that can evolve. Participants emphasized a nearer design horizon of 100 years as a meaningful unit for institutional and infrastructural commitments, while also using speculative devices to explore what happens as protocols outlive their creators.
@@ 4.2. Process and Setting
The workshop was structured as a relay across six parallel story worlds. Participants were divided into tables of five and worked through six sequential rounds, with each table contributing to every scenario in turn. After the sixth round, each story was returned to its original authors. This allowed them to see how their initial protocol had drifted, mutated, or been reinterpreted by others, rendering interpretive drift visible and allowing them to experience how protocols evolve when traversed across many hands with only partial context.
@@
Each story world began with a “Blue Build 1” phase, where a team proposed a knowledge-preservation protocol for the next 100 years. This required teams to articulate both a technical stack (e.g., encoding methods, storage substrates, verification technologies) and a social milieu (e.g., maintenance institutions, stewarding communities, norms, governance practices). In the subsequent round, a new team inherited this design and introduced a “Red Crisis 1” intervention that represented political, ecological, social, or technological shocks that destabilized or reinterpreted the original protocol. The third round, Blue Build 2, tasked yet another team with continuing the story under these altered conditions for the next 300 years. They had to adapt, improvise, or recompose the preservation protocol in light of the disruptions and ambiguities left by their predecessors. This was followed by Red Crisis 2, in which yet another group introduced further disruptions, forcing reinterpretations, repair, or workaround practices. This alternating pattern of “build” and “crisis” repeats one last time with Blue Build 3 extending the protocol into the far future (the next 600 years), after which Red Crisis 3 introduced a last set of shocks, before the story was returned to the original authors. Over the course of this process, each knowledge artifact developed a layered, multi-century “protocol biography”: a record of how initial rules, assumptions, and incentives accumulated unintended consequences over centuries.
@@
The workshop was conducted entirely through six shared Google Doc files, one per story world, edited sequentially by all thirty participants. Each table deliberated in person while one designated note-taker captured its decisions around technology stacks, social milieus, crises, and state conditions directly into the evolving document. Because teams worked from what previous groups had written – brief bullet points, shorthand, half-formed sentences, and ambiguous phrasing – later teams frequently had to infer missing rationale, decide whether a line was meant as a firm rule or a tentative thought, and interpret text that reflected the stylistic quirks of an earlier note-taker rather than any stable consensus. No communication was allowed between tables until the final debrief, amplifying the interpretive gaps and making drift a constitutive part of the design process.
This structure deliberately reproduced the dynamics of protocol evolution in the wild, where specifications, proposals, patches, and lived practices often sediment into infrastructures without any single authorial center or developmental trajectory. As in actual decentralized systems, protocols themselves were the design artifacts, rather than the fictional worlds, but the rules, mediations, and assumptions that structured their long-term behavior. The shared documents functioned as proto-infrastructures that were mutable, cumulative, and open for reinterpretation, reflecting how technical rules and social practices co-produce one another over time.
@@
Interpretation of the fictional outputs occurred at multiple stages. During the workshop, participants themselves engaged in real-time interpretation as they inherited and responded to prior teams’ designs—making sense of ambiguous handovers, inferring unstated rationales, and deciding how to extend or contest the emerging protocol worlds. In the post-workshop phase, the research team conducted follow-up interviews and collaborative analysis sessions with participants, during which both researchers and participants reflected on the scenarios, surfaced implicit assumptions, and identified recurring patterns such as drift, jam, and ossification. This layered interpretive process—participant sense-making during the relay, followed by researcher-participant collaborative analysis—grounds the methodological insights in both experiential and analytic registers.
@@ 4.3.2. Red Crisis One (2120–2125): Authenticity Under Siege
4.3.3. Blue Build Two and Red Crisis Two (2125–2425): Hash Wars and Environmental Entanglement
Subsequent teams attempted to repair the jam by introducing new layers of protocol: comet-based “carbon dating” of engraved replicas, sandboxed LLMs to reconstruct pre-LLM word distributions, and a blockchain-verified hash of the “oldest known Wikipedia.” Interviewed participants described this as an attempt to build a “protocol hardness”—a durability that seeks to anchor truth in mechanisms that could not easily be spoofed or reinterpreted. At the same time, the social milieu shifted. AI researchers, amateur astronomy clubs, and archivists formed a coalition around maintaining a “clean” pre-LLM dataset. This coalition reinforced the idea that some slices of knowledge must remain stable referents, even as everything else is up for reinterpretation.
@@ 4.3.3. Blue Build Two and Red Crisis Two (2125–2425): Hash Wars and Environmental Entanglement
By 2425, distorted information had become culturally normalized, while undistorted data was derided as naïve or heretical. In this stage, a protocol designed to verify authenticity through intensive hash comparison and space-based signaling became entangled with heavy compute demands, regulatory oversight, and adversarial exploitation, ultimately producing second-order effects such as environmental externalities, political clampdowns, and a cultural turn that celebrated distortion over fidelity. The interviews helped clarify this trajectory as a risk-aware framing of long-term preservation: participants explicitly discussed political, economic, social, legal, and environmental constraints as integral to protocol design. Crisis recovery was not just about restoring systems, but about maintaining legitimacy under shifting constraints.
@@ 4.4. Design Principles and Mediating Practices
Interpretation occurred at multiple moments across the relay. During the workshop, participants themselves interpreted and reworked inherited protocol fragments under constrained conditions. After the workshop, the research team conducted a second-order analytic pass, drawing on field notes, the evolving documents, and follow-up interviews to trace how protocol rules, mediating practices, and second-order dynamics emerged across rounds. Minimal validity checks included adversarial critique built into the Red Crisis phases, participant reflection during the final debrief, and post-hoc elicitation interviews that allowed participants to confirm, contest, or reinterpret the researchers’ analytic readings.
@@
Third, participants treated crisis and recovery as intrinsic to protocol design. Several emphasized the need for “recovery protocols” capable of restoring function after disruption without sacrificing interpretability. This anticipatory stance toward failure differentiates Protocol Futuring from more static forms of design fiction, which often portray artifacts as working as intended within a closed narrative world. Here, resilience, repair, and flexible adaptation were core design considerations.
Fourth, participants explicitly noted how philosophical and cultural attitudes—such as skepticism toward historical truth or tendencies toward artistic reinterpretation—might shape the long-term fate of preservation systems. Rather than assuming fidelity as a universal good, they saw value in designing infrastructures that honor historical authenticity while permitting creative reimagination. This orientation positions Protocol Futuring as a way to explore how future publics might negotiate the tension between accuracy and meaning.
Finally, challenges around collaboration and communication surfaced repeatedly through metaphors of “different islands.” Because teams worked asynchronously through textual handovers, misinterpretation and semantic drift were routine features of the exercise. Participants proposed concepts such as “lingua defense”—shared languages, standards, and interpretive conventions—as cultural protocols to mitigate these breakdowns. This insight reframes communication and consensus mechanisms as preservation infrastructures in their own right, revealing how much interpretive labor underpins even seemingly technical systems.
@@ 5.1. Temporality, Second-Order Dynamics, and Protocol “Memory”
In Knowledge Futurama, participants collectively designed archival protocols intended to persist for centuries. The resulting fictions centered not only on immediate trade-offs (e.g., privacy vs. access, centralization vs. federation) but also on how those choices sediment into long-term trajectories: protocols accumulating layers of patches, losing legibility, or being repurposed for unexpected ends. This echoes Ribes and Finholt (2007)’s long-term planning of infrastructures, where tensions between flexibility and stability, project timescales and infrastructural timescales, become central. Our speculative archive worlds dramatized, for instance, how a global archive protocol might gradually drift away from its founding rationale as institutional memory erodes and new stakeholders reinterpret its rules. The second-order dynamics surfaced in Knowledge Futurama should therefore be read as reflective hypotheses—patterns made thinkable through speculative practice and paired with candidate observables for future empirical follow-up, rather than as determinate claims about how archival infrastructures will evolve.
@@
Seen through STS work on temporality, these stories can be understood as miniature sociotechnical imaginaries attached to particular protocol regimes. Jasanoff and Kim (2015) argue that large-scale projects—from nuclear power to smart grids—are guided by collectively held imaginaries of national futures. In our method, participants effectively articulate competing imaginaries of protocolized futures: worlds where archival AIs act as guardians of planetary memory versus worlds where protocol-enforced scarcity creates exclusion and information black markets. Protocol Futuring thus surfaces how different imaginaries are encoded into protocol choices, and how those imaginaries, in turn, shape what second-order phenomena are even thinkable.
@@
Our scenarios also foreground protocol “memory”: the ways in which protocols remember (or fail to remember) the world that produced them. Bowker’s work on biodiversity databases shows how infrastructures embody particular classificatory decisions that persist long after their original context has faded (Bowker, 2000). Similarly, the archive stories exhibit situations where only a small priesthood remembers why certain safeguards exist by the year 2500. Participants proposed design responses—such as embedding narrative origin stories or mythologies into code comments, protocol documents, or ritual practices—that explicitly tie technical rules to their founding controversies. This points toward protocol design strategies that treat documentation, storytelling, and education as first-class components of long-term governance, not mere afterthoughts.
@@
Finally, Protocol Futuring helps articulate potential temporal failure modes of protocols as interpretive hypotheses for further inquiry. Through our fictions, we can distinguish candidate patterns such as: (a) drift, where practice may diverge from formal rules; (b) jam, where incompatible incentives or overloaded dependencies could lead to contention and deadlock; and (c) ossification, where the cost or risk of change may become so high that even beneficial modifications are repeatedly deferred. These second-order dynamics are not predictions but interpretive hypotheses generated through speculative practice—plausible trajectories that warrant empirical investigation. Each pattern maps onto potentially observable phenomena in real infrastructures (e.g., workarounds and shadow systems, prolonged standardization disputes, or entrenched legacy protocols). Future empirical work can focus on candidate observables such as divergence between protocol text and workaround artifacts (drift), frequency and distribution of change proposals (ossification), or patterns of stakeholder contestation (jam).